### v1

@@ IV-B1 Understanding telecom domain knowledge
A thorough understanding of the above telecom knowledge is not only vital for the development of new technologies and services, but also for ensuring that systems are interoperable and secure. The depth of understanding in telecom knowledge directly impacts the ability to innovate, secure, and solve problems within the telecom field. The integration of LLM models, trained with domain-specific datasets, offers promising avenues for automating knowledge generation and facilitating access to complex telecom content, thus bridging the gap between experts and general users.
@@
IV-B2 Training LLM models with telecom-specific data
Training LLM models with telecom-specific data involves curating and preprocessing vast amounts of domain-specific information to fine-tune the models, aiming to generate accurate and relevant content within the telecom field. This process is crucial as it tailors the LLM’s capabilities to understand and generate content that aligns with specific telecom requirements. It can be summarized by following steps:
@@ IV-B2 Training LLM models with telecom-specific data
  • Finally, it is worth noting that there are two main approaches to train LLM models, which are training the model from scratch or fine-tuning a general-domain LLM. In particular, training the model from scratch may produce better performance since the model can specialize in telecom language, but it is also time-consuming. On the other hand, the fine-tuning process adapts the pre-trained LLM to the telecom domain. This step involves training the model on the collected telecom-specific dataset, allowing it to adjust its parameters to better understand and generate telecom content. Fine-tuning enables the model to grasp the unique terminologies, concepts, and contexts of the telecom field, significantly enhancing its generation capabilities. Although fine-tuning a pre-trained LLM is much more efficient than training from scratch, the experiment in [100] proves that training the model on telecom-domain text from scratch can achieve better performance than fine-tuning a general-domain model.
@@ IV-B3 Using LLM to telecom knowledge-related generation tasks
After proper training or fine-tuning, using LLM models to generate telecom domain knowledge is a transformative approach that leverages the model’s ability to process and synthesize vast amounts of information into coherent, accessible content tailored to the needs of various stakeholders in the telecom field. This capability extends from generating summaries of complex technical documents to answering specific queries with detailed explanations, thereby facilitating a deeper understanding of telecom technologies, standards, and practices. In the following, we present some existing applications of telecom knowledge-related generation tasks.
Telecom-domain question answering: Question answering is one of the most well-known applications of LLM technologies. Using LLM models to answer domain-specific questions is grounded in the model’s ability to interpret and articulate complex information in a manner that is both comprehensive and understandable. For example, Soman et al. evaluated the capabilities and limitations of existing pre-trained general domain LLMs in [102], including GPT-3.5, GPT-4, Bard, and LLaMA. For instance, one telecom-domain question is ”What are the different 5G spectrum layers?” GPT-4 identifies the bands as below 1 GHz, 1-6 GHz and above 6 GHz, while LLaMA identifies the frequency bands as below 600 MHz, 600 MHz-24 GHz and above 24 GHz. These differences could be caused by different data sources of GPT-4 and LLaMA in the pre-training period. However, this could easily confuse or even mislead users without professional knowledge, which shows the importance of training a telecom-domain LLM specifically. Holm et al. [100] further investigate how various training methods can affect the model performance, e.g., pre-training a model using telecom knowledge from scratch or fine-tuning an existing general-domain model. In summary, LLM-enabled question answering democratizes access to advanced telecom knowledge, making it accessible to a broader audience, including researchers, practitioners, and the general public. In addition, LLM agents can also tailor the generated content based on the user’s level of expertise and specific interests. There is an increasing number of commercial LLM products for generative question answering over business documents, e.g., nexocode and Caryon. By leveraging the comprehensive understanding and generation capabilities of LLM technologies, the telecom industry can enhance the accessibility of complex information, support educational endeavours, and streamline development processes.
@@ IV-C2 LLM-aided code generation with multi-step scheduling
Previous sections have shown that LLM can be used for fundamental coding tasks. However, real-world telecom project development is usually much more complicated by including multi-step scheduling and several sub-tasks. Xiang et al. applied LLM to regenerate the code of existing studies in [18], and the authors suggested that ChatGPT does not respond well to monolithic prompts like ”implement this technique in the following steps”. Instead, a more practical method is to send a detailed modular prompt each time. Such a step-by-step approach is also investigated in [104] and [14]. Specifically, Mani et al. [104] applied LLM models to network graph manipulation, and the prompt design is decoupled into the application prompt and code generation prompt. Specifically, the application prompt can provide task-specific prompts based on templates and user queries, and then the code generation prompt can use plugins and libraries to instruct LLM models. The experiment shows that combining the LLM with proper libraries, such as GPT-4 and NetworkX, can achieve 88% and 78% coding accuracy for traffic analysis and network lifecycle management tasks, respectively. Du et al. investigated a more complicated coding task in [14] by using Verilog to build a Fast Fourier Transform (FFT) module. A failure is first observed by using the following prompt:
@@
Finally, we summarize some key lessons from existing studies on the use of LLM for code generation. Firstly, step-by-step prompt design is an important lesson that has been demonstrated in several existing studies[14, 18, 104]. Decoupling the complicated multi-step scheduling problem into several stages will lower the difficulty for LLM’s understanding. For instance, in 5G cloud RAN simulation, we can divide the network into cloud, edge, and users, and then use LLM to generate the code for each part sequentially. Secondly, examples and pseudo-code are important for code generation. The LLM has excellent ICL capabilities, quickly learning from examples and generalizing to other scenarios. Xiang et al. [18] also reveal that implementation with pseudocode first can produce stabilized data types and structures, avoiding other changes when implementing the following components. There have been many codes for the telecom field in GitHub and textbooks, taking advantage of these existing examples is crucial to use LLM techniques. Then, a significant amount of human effort can be saved in code generation by using LLM models for debugging and testing. Xiang et al. [18] also shows that most errors can be solved by sending the error message to the LLM. Many of these errors are related to data types, which can be avoided by specifying key variables’ data types. This lesson is also proved in [14], in which the LLM specified the data types of inputs and outputs to improve the reliability of existing code. Finally, LLM-aided coding can lower the requirement for professional knowledge [14, 18]. In particular, Du et al. [14] show that both undergraduate and graduate students can benefit from the assistance of LLM models, achieving comparable coding qualities. Xiang et al. [18] prove that undergraduate students can reproduce the results of some existing network studies by using the LLM.
@@ IV-D Network Configuration Generation
One notable development of LLM-aided network configuration is CloudEval-YAML [24], a benchmark that provides a realistic and scalable assessment framework specifically for YAML configurations in cloud-native applications. This benchmark utilizes a hand-crafted dataset and an efficient evaluation platform to thoroughly examine the performance of LLM models within this context. Dzeparoska et al. [107] have introduced a pioneering method that employs the few-shot learning capabilities of the LLM to automate the translation of high-level user intents into executable policies. This approach facilitates dynamic, automated management of applications without the necessity for predefined procedural steps. In a related vein, Wang et al. [108] have developed NETBUDDY, a multi-stage pipeline that leverages LLM models to translate high-level network policies specified in natural language into low-level device configurations. NETBUDDY first uses an LLM to convert the input into a formal specification, such as a data structure to express reachability. It then generates forwarding information and configuration scripts from the formal specification. Finally, NETBUDDY interacts with an LLM multiple times to sequentially provide topology, addressing details and prototype programs to automatically generate vendor-agnostic configurations for the switches and routers. The evaluation of the network emulator demonstrates NETBUDDY’s ability to enforce path policies and dynamically modify existing deployments. In addition, Mondal et al. [109] presented Verified Prompt Programming (VPP) to improve GPT-4’s ability to synthesize router configurations. VPP combines GPT-4 with verifiers like Batfish [120], which check configurations for syntax errors and semantic differences. Experiments showed that VPP presented 10×10$\times$10 × leverage performance for translating a Cisco configuration to Juniper format by identifying and fixing syntax errors, structural mismatches, attribute differences, and policy behaviours through 20 automated prompts. Implementing no-transit policies across 6 routers achieved 6×6$\times$6 × leverage performance with 12 automated prompts guiding GPT-4 to resolve syntax, topology, and semantic policy errors.
@@
Domain knowledge generation Telecom-domain question answering Question answering is the most well-known application of LLMs. It represents a significant step forward in the ongoing effort to bridge the knowledge gap in telecom and empower individuals and organizations within the field, including telecom question answering, literature summary and review, etc. General domain LLM models can also answer telecom-related questions, but fine-tuning a telecom-specific LLM can provide more reliable and professional answers. CoT prompting may improve the answer quality. The use of LLM techniques signifies a shift towards more efficient and accessible knowledge dissemination methods than any existing textbooks, websites, and tutorials for their comprehension and reasoning capabilities. 1) Building a telecom-domain LLM is a promising direction to make telecom-knowledge more accessible for both professional researchers and the public. 2) Automated troubleshooting is another promising application to automate the problem-solving process in telecom fields. 3) The LLM also have the potential of generating other language-related tasks, e.g., specifications and protocols.
@@
Network configuration generation Automatic network configuration generation by using LLM models. Using the LLM to generate network configurations automatically, and then verify by LLM models or humans. The prompt must be carefully designed due to the complexity of network configurations, e.g., dividing the prompts into the task-specific part and code generation part. The LLM enables efficient generation of network configurations, reducing manual effort and cost in the telecom industry. Applications include automatic network provisioning, performance tuning, security and compliance configuration, etc.
@@ IV-E Discussions and Analyses
Firstly, prompt design is critical for multi-step network generation problems. Telecom networks are complicated systems, and many tasks require dedicated design with multi-step scheduling. For example, the wireless code generation problem in Section IV-C2 proves that using a one-step prompt to generate the whole project is unpractical, e.g., generating a 64-point FFT module [14], and reproducing the results of existing publications [18]. Therefore, proper prompt design is critical to obtain the desired output. For instance, the prompt design [104] is decoupled into the application prompt and code generation prompt, in which the application prompt focuses on task-specific requirements, and the code generation prompt uses plugins and libraries to instruct LLM models.
Meanwhile, LLM-enabled generation can significantly save humane efforts. Existing studies have shown that LLM has excellent capabilities for code refactoring and validation, which are usually solved manually with considerable human effort. Applying such a technique to the telecom field will significantly save human labour on projecting coding, validating, and debugging. For instance, Zhang et al. [105] introduce that the LLM can successfully repair 86.71% programs for introductory level Python projects, and adding few-shot examples will raise the ratio to 96.50%. In addition, LLM models can also abstract fundamental knowledge in the network field from textbooks, journals, and specification standards, which avoids the time-consuming literature review process.
LLM models have been trained on many real-world datasets from web pages like Wikipedia, and they can be further fine-tuned on domain-specific datasets, e.g., telecom[110] and cybersecurity[121]. Despite the satisfactory performance, there is no guarantee for the correctness of the generated output. Such risks are avoided when the generated code or network configuration can be verified by pre-designed test cases. However, when using the LLM to summarize or extract knowledge from existing literature, the quality of generated knowledge is hard to validate. For example, LLM models may produce wrong numbers or units, and these mistakes can easily mislead users without professional knowledge. To this end, efficient validation schemes are crucial to evaluate the performance of generated solutions, especially for coding and network configuration problems. Human verification is a simplistic and straightforward approach, but it requires considerable human labour and can be inefficient. Therefore, automatic validation is the key to improve the overall efficiency of the whole pipeline, e.g., sending the code implementation error message to a LLM for automatic debugging[18], and using LLMs to validate the network configuration files.
@@ V LLM-enabled Classification Problems
[124] Building a security-specific LLM from scratch designed for detecting network cyber threats, involving several steps: data preparation, data tokenization, model training, and model deployment. SecurirtBERT showcases the powerful predictive capabilities of security-specific LLM models in identifying various types of attacks, significantly outperforming the traditional ML and DL models.
@@
[13] Fine-tuning several LLM models, e.g., BERT, distilled BERT, RoBERTa and GPT-2, to the telecom domain languages, and using them for 3GPP standard classification problems. With proper pre-processing and fine-tuning, the experiment in [13] can achieve an 80% accuracy even if only 20% of the text segments are used.
@@
V-A Motivations and Classification Capabilities of LLM Models
@@
Firstly, LLM technologies can contribute to telecom security by automated security language understanding and classification. Inspired by numerous advancements in NLP, LLM excels at explaining textual contents and transforming them into informative representations, such as GPT [134] and BERT [44]. With their strong capabilities, LLMs have recently shown exceptional superiority in attack detection, aiming to enhance the security of telecom networks[124]. Through fine-tuning pre-trained models or developing LLMs from scratch, LLMs retain the functional capabilities in general English while gaining a thorough understanding of the specialized security language, allowing LLM models to effectively identify and respond to security threats in telecom networks.
@@
In addition, LLM models can extract visual features from the dynamic and complex telecom environment. The integration of computer vision and image processing into the telecom field, such as equipping BSs with cameras to pinpoint user locations, can boost network efficiency in the dynamically changing wireless environment. Although primarily focusing on processing and understanding textual information, some LLM models also have remarkable image processing capabilities in vision-related tasks[136], including image-to-text generation [137] and object detection [138], etc. Consequently, this integration enables the LLM to analyze both visual and network data, which can effectively bridge the gap between textual and visual data analysis, leading to a more comprehensive approach for network management.

### v2

@@ IV-B1 Understanding telecom domain knowledge
A thorough understanding of the above telecom knowledge is not only vital for the development of new technologies and services, but also for ensuring that systems are interoperable and secure. The depth of understanding in telecom knowledge directly impacts the ability to innovate, secure, and solve problems within the telecom field. The integration of LLMs, trained with domain-specific datasets, offers promising avenues for automating knowledge generation and facilitating access to complex telecom content, thus bridging the gap between experts and general users.
@@
IV-B2 Training LLMs with telecom-specific data
Training LLMs with telecom-specific data involves curating and preprocessing vast amounts of domain-specific information to fine-tune the models, aiming to generate accurate and relevant content within the telecom field. This process is crucial as it tailors the LLM’s capabilities to understand and generate content that aligns with specific telecom requirements. It can be summarized by following steps:
@@ IV-B2 Training LLMs with telecom-specific data
  • Finally, it is worth noting that there are two main approaches to train LLMs, which are training the model from scratch or fine-tuning a general-domain LLM. In particular, training the model from scratch may produce better performance since the model can specialize in telecom language, but it is also time-consuming. On the other hand, the fine-tuning process adapts the pre-trained LLM to the telecom domain. This step involves training the model on the collected telecom-specific dataset, allowing it to adjust its parameters to better understand and generate telecom content. Fine-tuning enables the model to grasp the unique terminologies, concepts, and contexts of the telecom field, significantly enhancing its generation capabilities. Although fine-tuning a pre-trained LLM is much more efficient than training from scratch, the experiment in [115] proves that training the model on telecom-domain text from scratch can achieve better performance than fine-tuning a general-domain model.
@@ IV-B3 Using LLM to telecom knowledge-related generation tasks
After proper training or fine-tuning, using LLMs to generate telecom domain knowledge is a transformative approach that leverages the model’s ability to process and synthesize vast amounts of information into coherent, accessible content tailored to the needs of various stakeholders in the telecom field. This capability extends from generating summaries of complex technical documents to answering specific queries with detailed explanations, thereby facilitating a deeper understanding of telecom technologies, standards, and practices. In the following, we present some existing applications of telecom knowledge-related generation tasks.
Telecom-domain question answering: Question answering is one of the most well-known applications of LLM technologies. Using LLMs to answer domain-specific questions is grounded in the model’s ability to interpret and articulate complex information in a manner that is both comprehensive and understandable. For example, Soman et al. evaluated the capabilities and limitations of existing pre-trained general domain LLMs in [117], including GPT-3.5, GPT-4, Bard, and LLaMA. For instance, one telecom-domain question is ”What are the different 5G spectrum layers?” GPT-4 identifies the bands as below 1 GHz, 1-6 GHz and above 6 GHz, while LLaMA identifies the frequency bands as below 600 MHz, 600 MHz-24 GHz and above 24 GHz. These differences could be caused by different data sources of GPT-4 and LLaMA in the pre-training period. However, this could easily confuse or even mislead users without professional knowledge, which shows the importance of training a telecom-domain LLM specifically. Holm et al. [115] further investigate how various training methods can affect the model performance, e.g., pre-training a model using telecom knowledge from scratch or fine-tuning an existing general-domain model. In summary, LLM-enabled question answering democratizes access to advanced telecom knowledge, making it accessible to a broader audience, including researchers, practitioners, and the general public. In addition, LLM agents can also tailor the generated content based on the user’s level of expertise and specific interests. There is an increasing number of commercial LLM products for generative question answering over business documents, e.g., nexocode and Caryon. By leveraging the comprehensive understanding and generation capabilities of LLM technologies, the telecom industry can enhance the accessibility of complex information, support educational endeavours, and streamline development processes.
@@ IV-C2 LLM-aided code generation with multi-step scheduling
Previous sections have shown that LLM can be used for fundamental coding tasks. However, real-world telecom project development is usually much more complicated by including multi-step scheduling and several sub-tasks. Xiang et al. applied LLM to regenerate the code of existing studies in [18], and the authors suggested that ChatGPT does not respond well to monolithic prompts like ”implement this technique in the following steps”. Instead, a more practical method is to send a detailed modular prompt each time. Such a step-by-step approach is also investigated in [119] and [14]. Specifically, Mani et al. [119] applied LLMs to network graph manipulation, and the prompt design is decoupled into the application prompt and code generation prompt. Specifically, the application prompt can provide task-specific prompts based on templates and user queries, and then the code generation prompt can use plugins and libraries to instruct LLMs. The experiment shows that combining the LLM with proper libraries, such as GPT-4 and NetworkX, can achieve 88% and 78% coding accuracy for traffic analysis and network lifecycle management tasks, respectively. Du et al. investigated a more complicated coding task in [14] by using Verilog to build a Fast Fourier Transform (FFT) module. A failure is first observed by using the following prompt:
@@
Finally, we summarize some key lessons from existing studies on the use of LLM for code generation. Firstly, step-by-step prompt design is an important lesson that has been demonstrated in several existing studies[14, 18, 119]. Decoupling the complicated multi-step scheduling problem into several stages will lower the difficulty for LLM’s understanding. For instance, in 5G cloud RAN simulation, we can divide the network into cloud, edge, and users, and then use LLM to generate the code for each part sequentially. Secondly, examples and pseudo-code are important for code generation. The LLM has excellent ICL capabilities, quickly learning from examples and generalizing to other scenarios. Xiang et al. [18] also reveal that implementation with pseudocode first can produce stabilized data types and structures, avoiding other changes when implementing the following components. There have been many codes for the telecom field in GitHub and textbooks, taking advantage of these existing examples is crucial to use LLM techniques. Then, a significant amount of human effort can be saved in code generation by using LLMs for debugging and testing. Xiang et al. [18] also shows that most errors can be solved by sending the error message to the LLM. Many of these errors are related to data types, which can be avoided by specifying key variables’ data types. This lesson is also proved in [14], in which the LLM specified the data types of inputs and outputs to improve the reliability of existing code. Finally, LLM-aided coding can lower the requirement for professional knowledge [14, 18]. In particular, Du et al. [14] show that both undergraduate and graduate students can benefit from the assistance of LLMs, achieving comparable coding qualities. Xiang et al. [18] prove that undergraduate students can reproduce the results of some existing network studies by using the LLM.
@@ IV-D Network Configuration Generation
One notable development of LLM-aided network configuration is CloudEval-YAML [24], a benchmark that provides a realistic and scalable assessment framework specifically for YAML configurations in cloud-native applications. This benchmark utilizes a hand-crafted dataset and an efficient evaluation platform to thoroughly examine the performance of LLMs within this context. Dzeparoska et al. [122] have introduced a pioneering method that employs the few-shot learning capabilities of the LLM to automate the translation of high-level user intents into executable policies. This approach facilitates dynamic, automated management of applications without the necessity for predefined procedural steps. In a related vein, Wang et al. [123] have developed NETBUDDY, a multi-stage pipeline that leverages LLMs to translate high-level network policies specified in natural language into low-level device configurations. NETBUDDY first uses an LLM to convert the input into a formal specification, such as a data structure to express reachability. It then generates forwarding information and configuration scripts from the formal specification. Finally, NETBUDDY interacts with an LLM multiple times to sequentially provide topology, addressing details and prototype programs to automatically generate vendor-agnostic configurations for the switches and routers. The evaluation of the network emulator demonstrates NETBUDDY’s ability to enforce path policies and dynamically modify existing deployments. In addition, Mondal et al. [124] presented Verified Prompt Programming (VPP) to improve GPT-4’s ability to synthesize router configurations. VPP combines GPT-4 with verifiers like Batfish [135], which check configurations for syntax errors and semantic differences. Experiments showed that VPP presented 10×10$\times$10 × leverage performance for translating a Cisco configuration to Juniper format by identifying and fixing syntax errors, structural mismatches, attribute differences, and policy behaviours through 20 automated prompts. Implementing no-transit policies across 6 routers achieved 6×6$\times$6 × leverage performance with 12 automated prompts guiding GPT-4 to resolve syntax, topology, and semantic policy errors.
@@
Domain knowledge generation Telecom-domain question answering Question answering is the most well-known application of LLMs. It represents a significant step forward in the ongoing effort to bridge the knowledge gap in telecom and empower individuals and organizations within the field, including telecom question answering, literature summary and review, etc. General domain LLMs can also answer telecom-related questions, but fine-tuning a telecom-specific LLM can provide more reliable and professional answers. CoT prompting may improve the answer quality. The use of LLM techniques signifies a shift towards more efficient and accessible knowledge dissemination methods than any existing textbooks, websites, and tutorials for their comprehension and reasoning capabilities. 1) Building a telecom-domain LLM is a promising direction to make telecom-knowledge more accessible for both professional researchers and the public. 2) Automated troubleshooting is another promising application to automate the problem-solving process in telecom fields. 3) The LLM also have the potential of generating other language-related tasks, e.g., specifications and protocols.
@@
Network configuration generation Automatic network configuration generation by using LLMs. Using the LLM to generate network configurations automatically, and then verify by LLMs or humans. The prompt must be carefully designed due to the complexity of network configurations, e.g., dividing the prompts into the task-specific part and code generation part. The LLM enables efficient generation of network configurations, reducing manual effort and cost in the telecom industry. Applications include automatic network provisioning, performance tuning, security and compliance configuration, etc.
@@ IV-E Discussions and Analyses
Firstly, multi-step planning capabilities are crucial for telecom-related generation tasks. Telecom networks are large-scale complicated systems, and many tasks require dedicated planning and scheduling. For example, the study in [14] demonstrated that using a one-step prompt to generate a complicated 64-point FFT module is impractical, while step-by-step planning can achieve a satisfactory result. Similarly, step-by-step reasoning and planning are also useful to reproduce the results of existing publications for code generation problems [18]. Therefore, multi-step planning, i.e., step-by-step prompt design, is critical to obtain the desired output. For instance, the prompt design [119] is decoupled into the application prompt and code generation prompt, in which the application prompt focuses on task-specific requirements, and the code generation prompt uses plugins and libraries to instruct LLMs.
Meanwhile, LLM-enabled generation can significantly save humane efforts. Existing studies have shown that LLM has excellent capabilities for code refactoring and validation, which are usually solved manually with considerable human effort. Applying such a technique to the telecom field will significantly save human labour on projecting coding, validating, and debugging. For instance, Zhang et al. [120] introduce that the LLM can successfully repair 86.71% programs for introductory level Python projects, and adding few-shot examples will raise the ratio to 96.50%. In addition, LLMs can also abstract fundamental knowledge in the network field from textbooks, journals, and specification standards, which avoids the time-consuming literature review process.
LLMs have been trained on many real-world datasets from web pages like Wikipedia, and they can be further fine-tuned on domain-specific datasets, e.g., telecom[125] and cybersecurity[136]. Despite the satisfactory performance, there is no guarantee for the correctness of the generated output. Such risks are avoided when the generated code or network configuration can be verified by pre-designed test cases. However, when using the LLM to summarize or extract knowledge from existing literature, the quality of generated knowledge is hard to validate. For example, LLMs may produce wrong numbers or units, and these mistakes can easily mislead users without professional knowledge. To this end, efficient validation schemes are crucial to evaluate the performance of generated solutions, especially for coding and network configuration problems. Human verification is a simplistic and straightforward approach, but it requires considerable human labour and can be inefficient. Therefore, automatic validation is the key to improve the overall efficiency of the whole pipeline, e.g., sending the code implementation error message to a LLM for automatic debugging[18], and using LLMs to validate the network configuration files.
@@ V LLM-enabled Classification Problems
[139] Building a security-specific LLM from scratch designed for detecting network cyber threats, involving several steps: data preparation, data tokenization, model training, and model deployment. SecurirtBERT showcases the powerful predictive capabilities of security-specific LLMs in identifying various types of attacks, significantly outperforming the traditional ML and DL models.
@@
[13] Fine-tuning several LLMs, e.g., BERT, distilled BERT, RoBERTa and GPT-2, to the telecom domain languages, and using them for 3GPP standard classification problems. With proper pre-processing and fine-tuning, the experiment in [13] can achieve an 80% accuracy even if only 20% of the text segments are used.
@@
V-A Motivations and Classification Capabilities of LLMs
@@
Firstly, LLM technologies can contribute to telecom security by automated security language understanding and classification. Inspired by numerous advancements in NLP, LLM excels at explaining textual contents and transforming them into informative representations, such as GPT [149] and BERT [53]. With their strong capabilities, LLMs have recently shown exceptional superiority in attack detection, aiming to enhance the security of telecom networks[139]. Through fine-tuning pre-trained models or developing LLMs from scratch, LLMs retain the functional capabilities in general English while gaining a thorough understanding of the specialized security language, allowing LLMs to effectively identify and respond to security threats in telecom networks.
@@
In addition, LLMs can extract visual features from the dynamic and complex telecom environment. The integration of computer vision and image processing into the telecom field, such as equipping BSs with cameras to pinpoint user locations, can boost network efficiency in the dynamically changing wireless environment. Although primarily focusing on processing and understanding textual information, some LLMs also have remarkable image processing capabilities in vision-related tasks[151], including image-to-text generation [152] and object detection [153], etc. Consequently, this integration enables the LLM to analyze both visual and network data, which can effectively bridge the gap between textual and visual data analysis, leading to a more comprehensive approach for network management.